find the inverse of f(x) = -x-4
Answer by stanbon(75887)

find the inverse of f(x) = -x-4
y = -x-4
--
1st: Interchange x and y.
x = -y-4
2nd: Solve for "y":
y = -x-4
That is the inverse.
=======================
Cheers,
Stan H.